Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
gherardini - 23:02 Wednesday 31 May 2023 (60391) Print this report
Operator Report - Morning shift

This afternoon shift was dedicated to: electrostatic charge measure on WI and NE and Q factor measurement, the work was carried out by Paolo with the ITF locked in LOW_NOISE_2 state, activity completed at 19:30UTC; then after ~1 hour of lock I manually unlocked the ITF to start FSR scan (from 19:40UTC to 20:40UTC), during this period the west green unlocked from time to time, I tried to restart the scan as faster then possible;
I left the ITF locked in LOW_NOISE_2 with the failsafe engaged.

Sub-system reports

Air Conditioning
at 16:50UTC I switched the central building clean room air conditioning system to reduced flux.

Images attached to this report
Comments to this report:
soldani - 16:27 Thursday 01 June 2023 (60408) Print this report
CEB Clean rooms AHU is off (with cold & warm water valves closed) since today @16.00 local time
tournefi - 18:27 Thursday 01 June 2023 (60414) Print this report

These are the FSR scans done yesterday. As mentioned by Fabio these were interrupted from time to time. Could experts check what was the problem with the green?

Images attached to this comment
gosselin, derossi - 18:35 Thursday 01 June 2023 (60413) Print this report

We took a look at the unlocks of the green during the scans.
They seem to be triggered by an oscillation at around 56 kHz that suddenly appears on the error signal (see figure 1)
By taking a look at the spectrum and the fft time of the unlocks we can spot several lines (figure 2 and figure 3).
The 70 kHz line is the phase modualtion used to lock the arms.
There are two fixed lines at 13.8 kHz and 27.6 kHz. They might be generated by the beating between the 70 kHz and the 56 kHz.... or the countrary. 
We can also spot a line at around 6 kHz which is increasing in frequency. It seems that there is a beating between this line and 50 kHz (see figure 4)

We do not have clear explanations for those oscillations yet (fixed and wandering), could they be due to some mechanical mode of the WE mirror?

We quickly look an other unlock which shows a similar but different behavior (figure 5)

Images attached to this comment
Search Help